OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (10, 11) m = -9

OpenStudy (anonymous):

y = -9x +101

OpenStudy (anonymous):

y = mx + b. let's plug in m and x and y from above: 11 = -9(10) + b 11 = -90 + b b = 101 Then: y = -9x + 101

OpenStudy (anonymous):

Give the equation y = mx + b, since you know the value of the slope (m = -9) and a coordinate on the line (10, 11), you can substitute these numbers into the equation giving you 11 = -9*10 + b, then solve for b.